# Large-Scale Calculations of $\beta$-Decay Rates and Implications for $r$-Process Nucleosynthesis

### v1

@@ 4 Conclusions
In this Letter, we presented the state-of-the-art global predictions of $\beta$-decay rates based on the relativistic nuclear EDF theory with DD-PC1 and DD-PCX functionals, including the effect of axial deformation of nuclei. The new rates agree well with the deformed non-relativistic SkO’ EDF predictions (Ney et al., 2020). These $\beta$-decay rates suggest slower $\beta$ decays, especially past the $N=126$ neutron shell closure, and are distinctly different from the phenomenological FRDM (Möller et al., 2019) or spherical relativistic EDF D3C* (Marketin et al., 2016) tabulations, which have been commonly used in $r$-process simulations. This significantly reduces the flow of neutron-induced fission, which affects the final abundance patterns. In less neutron-rich environments, such as magneto-rotational supernovae, the $\beta$-decay rates may compete with the expansion timescale of the ejecta, thereby affecting heavy nucleosynthesis. Since most astrophysical analyses depend on specific global $\beta$-decay models, such a choice may affect our understanding of the astrophysical environment of the $r$-process. The results presented in this work constitute a step in our effort towards comprehensive theoretical predictions based on the quantified nuclear EDF framework, which will enable a controlled assessment of uncertainties due to different nuclear properties.
5 Acknowledgments
@@
This work was supported by the U.S. Department of Energy under Award No. DOE-DE-NA0004074 (NNSA, the Stewardship Science Academic Alliances program), the DOE Office of Science under Grants DE-SC0013365 and DE-SC0023175 (Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing), and by the U.S. National Science Foundation under grant number 21-16686 (NP3M). Computational resources were provided in part by the Institute for Cyber-Enabled Research at Michigan State University. This work used TAMU ACES at TAMU through allocation PHY250120 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants No. 2138259, 2138286, 2138307, 2137603, and 2138296.

### v2

@@ IV Conclusions
In this paper, we presented the state-of-the-art global predictions of $\beta$-decay rates, including allowed and first-forbidden transitions, based on the relativistic nuclear EDF theory with DD-PC1 and DD-PCX functionals, including the effect of axial deformation of nuclei. The new rates agree well with the deformed non-relativistic SkO’ EDF predictions [Ney2020]. These $\beta$-decay rates suggest slower $\beta$ decays, especially past the $N=126$ neutron shell closure, and are distinctly different from the phenomenological FRDM [Moller2019] or spherical relativistic EDF D3C* [Marketin_RQRPA_2016] tabulations, which have been commonly used in $r$-process simulations. This significantly reduces the flow of neutron-induced fission, which affects the final abundance patterns. In less neutron-rich environments, such as magneto-rotational supernovae, the $\beta$-decay rates may compete with the expansion timescale of the ejecta, thereby affecting heavy nucleosynthesis. Since most astrophysical analyses depend on specific global $\beta$-decay models, such a choice may affect our understanding of the astrophysical environment of the $r$-process. The results presented in this work constitute a step in our effort towards comprehensive theoretical predictions based on the quantified nuclear EDF framework, which will enable a controlled assessment of uncertainties due to different nuclear properties.
V Acknowledgments
@@
We are grateful to Kyle Godbey for developing a website (https://rpx.ascsn.net) to display our results interactively. This work was supported by the U.S. Department of Energy under Award No. DOE-DE-NA0004074 (NNSA, the Stewardship Science Academic Alliances program), the DOE Office of Science under Grants DE-SC0013365 and DE-SC0023175 (Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing), and by the U.S. National Science Foundation under grant number 21-16686 (NP3M). Computational resources were provided in part by the Institute for Cyber-Enabled Research at Michigan State University. This work used TAMU ACES at TAMU through allocation PHY250120 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants No. 2138259, 2138286, 2138307, 2137603, and 2138296.